OpenStudy (lisa123):

Find all solutions in the interval [0, 2π).

OpenStudy (lisa123):

cos2x + 2 cos x + 1 = 0

OpenStudy (lisa123):

Pleaseeeee Help me I don't understand!

OpenStudy (mathmate):

Is the question A. \(cos2x + 2 cos x + 1 = 0\) or B. \(cos^2x + 2 cos x + 1 = 0\)

OpenStudy (lisa123):

The second one. (B)

OpenStudy (anonymous):

general rule for quadratic form

OpenStudy (anonymous):

(cos+1)^2=0

OpenStudy (anonymous):

cosx=-1

OpenStudy (lisa123):

ok so do I replace cosx with -1 in the equation?

OpenStudy (mathmate):

@Catch.me means that the solution for cos(x) is cos(x)=-1. Now you have to solve for x from the standard graph of cosine(x) between 0 and 2pi.  Here's what the graph looks like, from which you can interpret the value of x when cos(x)=-1.

OpenStudy (lisa123):

So if cos x=-1 x =3pi/2?

OpenStudy (mathmate):

 Not really, but hope this will make it more clear.